OpenStudy (nuccioreggie):

The regular price of a pair of jeans is $49.50. During a sale, the jeans were marked as 15% off. What was the approximate price of the jeans during the sale? $3.30 $7.43 $42.08 $56.93

OpenStudy (nuccioreggie):

@LegendarySadist

OpenStudy (nuccioreggie):

@SickTwistd

OpenStudy (anonymous):

7.43

OpenStudy (anonymous):

49.50 x .15

OpenStudy (nuccioreggie):

No ur wrong its 49.50÷15=3.30

OpenStudy (anonymous):

No you have to multiply to get the percent of something

OpenStudy (nuccioreggie):

@SickTwistd

OpenStudy (anonymous):

turn it into a decimal and multiply 15% would be .15

OpenStudy (anonymous):

LOL ULTIMATE BATTLE OF DEATH

OpenStudy (triciaal):

full price = 100% 15% off means you pay (100-15) = 85% of the full price 85/100*49 =

OpenStudy (tkhunny):

What is 15% of $49.50. This is the Discount. $49.50 - Discount = Sale Price.

OpenStudy (triciaal):

you could also do 49*15% and then subtract from 49

OpenStudy (nuccioreggie):

41.65

OpenStudy (nuccioreggie):

I got 7.35

OpenStudy (anonymous):

ugh

OpenStudy (nuccioreggie):

I got 7.43 i s that correct

OpenStudy (anonymous):

Told you. I was trying to help you.

OpenStudy (tkhunny):

49.50 * 0.15 = 7.43 -- You have found the discount. You have not found the Sale Price.

OpenStudy (nuccioreggie):

What the answer is C

OpenStudy (nuccioreggie):

Heloooooooo

OpenStudy (tkhunny):

Subtract and see. What do you get?

OpenStudy (nuccioreggie):

I get C

OpenStudy (nuccioreggie):

@tkhunny. I got C

OpenStudy (tkhunny):

Well, there you go. Move on to the next one with a little more confidence.